OpenStudy (precal):

Use the given information about f ' and f '' to find f(x) f " (x)=2 f '(2)=5 f(2)=10

OpenStudy (ikram002p):

integrate ^^

rvc (rvc):

correct sis

OpenStudy (precal):

integrate f ' (2) ?

OpenStudy (ikram002p):

Hint :- \(f'(x)=\int f''(x) \)

rvc (rvc):

yes

OpenStudy (ikram002p):

nope not f'(2) start with integrate f''(x)=2 do you know how to do that?

OpenStudy (precal):

ok so f(x)=2x correct?

OpenStudy (precal):

that can't be it because f(2) would not be 10

OpenStudy (ikram002p):

nope , check the hint above \(f'(x) = \int f''(x) +c \)

OpenStudy (precal):

I really don't know how to do these problems. I do better with functions.

OpenStudy (ikram002p):

so u have \(f'(x) = 2x+c\) since you know f'(2)=5 try to find c

OpenStudy (ikram002p):

ohkk :) only u are agree to this ?

rvc (rvc):

5=10+c

OpenStudy (precal):

ok I wil try

OpenStudy (precal):

c=1 ?

OpenStudy (ikram002p):

haha not try lol i wont continue unless ur agree to that ,so shall we start over ?

OpenStudy (precal):

yes

OpenStudy (ikram002p):

good :D so if i asked you whats the derivative of f(x) what do you say ?

OpenStudy (precal):

I truly am not interested in the solution. I really am going for learning how to do the problem so that I can do other similar problems and do well on my exam

OpenStudy (precal):

derivative of 2x is 2

OpenStudy (ikram002p):

no i meant any f(x) , u say derevative of f(x) is f'(x) ok ?

OpenStudy (ikram002p):

well , i would like to teach you :D

OpenStudy (precal):

ok I am sorry I thought we had established that it was 2x but yes

OpenStudy (precal):

yes the derivative of f(x) is f ' (x) the derivative of f '(x) is f " (x)

OpenStudy (ikram002p):

good :) now you know that integrate is the inverse of derevative right ?

OpenStudy (precal):

yes also, known as the antiderivative

OpenStudy (ikram002p):

cool :D

OpenStudy (ikram002p):

now what is the integration of f''(x) ? since we know that derevative f'(x) is f''(x) ?

OpenStudy (precal):

f ' (x)

OpenStudy (ikram002p):

yep your almost correct , we need c constant to add since (f'(x)+c)=f''(x) + c' and since c is constant then c' is zero Hence :- (f'(x)+c)=f''(x) so we are not sure if integrate f''(x) is exactly f'(x) it might have a constant addition agree to this ?

OpenStudy (ikram002p):

so we say \(\int f''(x) = ?\)

OpenStudy (precal):

need a moment to think

OpenStudy (ikram002p):

ok take ur time :D

OpenStudy (precal):

sorry had to chase off a woodpecker (he was on my house)

OpenStudy (precal):

but the integral of f "(x) should just be f '(x) + c as you stated earlier

OpenStudy (ikram002p):

lol i like that bird :P ur correct , so understood ?

OpenStudy (precal):

yes but still not sure what to do ok I will write what I think I should do and you can correct me. I am sure I am incorrect

OpenStudy (ikram002p):

ok ^^

OpenStudy (precal):

f(2)=10 is just a point (2,10) this just gives me a location f ' (2) =5 means that when I take the first derivative and sub 2 I should get 5 and this tells me that the function is increasing since 5 is a positive number or f '(2)>0 f"(x)=2 tells me that the function is concave up

OpenStudy (precal):

ok am I trying to find the function itself using the two values given f(2)=10 and f'(2)=5? I am not even sure I understand the directions

OpenStudy (ikram002p):

well check this \(f'(x) =\int f''(x) +C \) right ? then \(f'(x) =\int 2 dx +C \) can you find f'(X) now with respect to c ?

OpenStudy (precal):

f '(2)=2x +c 5=2(2)+c 5=4+c 1=c so 2x+c 2x+1 so I just need to integrate that and that is my f(x) correct?

OpenStudy (precal):

x^2 +x is my original function ok let me check all of the given information x=2 2^2+2 =4+2=6 so f(2)=10 does not work out guess I am still not sure

OpenStudy (ikram002p):

ok for f'(x) ur correct f'(x)=2x+1 :D now the same idea with f(x) dnt forget to add c \(f(x)=\int f'(x) +c\)

OpenStudy (precal):

ok then let me write on my paper and think this through

OpenStudy (ikram002p):

ohkkk :)

OpenStudy (precal):

ok my function is x^2+x+4 so let me check if x = 2 2^2+2+4=4+2=4=10 so f(2) = 10 f ' (x) = 2x+1 if f"(2)=2(2)+1=5 so that checks out as well

OpenStudy (ikram002p):

you are correct !

OpenStudy (ikram002p):

^^ well done

OpenStudy (precal):

thanks I have another one to tackle but I think I will do it later, gotta go run errands and chase off my woodpecker. I really do have an annoying bird attacking my house. btw Thanks for helping me understand this problem better.

OpenStudy (ikram002p):

np :) ur wlc PS :- don't hurt that bird ok :P

OpenStudy (precal):

no just chasing it off. They always go to the other houses that are made of wood. Mine is not wood, just fake wood. Once they realize this, they leave anyways.

OpenStudy (ikram002p):

huh ok :P good luck !